A paetical moving in xy plane such that its velocity =5i+4j find eqation of trajectory of partical if starts to t=0from rest of origin

Grade:12th pass

1 Answers

Mahima Kanawat
6 years ago
X={ucos©]t = 5tY={usin©}t~I/2gt^2X=5t ......(t=x/5)Y=4t~1/2gt^2Y=4x/5~1/2g(x/5)^2On solving5Y=4X~X^2